OpenStudy (anonymous):

Translate the following situation into an equation. Do not solve. “Suppose a crate of oranges weighs eight pounds more than a crate of bananas and a crate of kiwi weighs one third as much as a crate of bananas. Together the three crates weigh 43 pounds. What is the weight of the crate of bananas?

OpenStudy (anonymous):

simultaneous equations: (1) o = b + 8 (2) k/3 = b (3) o + b + k = 43